OpenStudy (milena):

manufacturer can produce digital recorders at a cost of $50 apiece. It is estimated that if the recorders are sold for p dollars apiece, consumers will buy q 120 p recorders each month. Express the manufacturer’s profit P as a func- tion of q. b. What is the average rate of profit obtained as the level of production increases from q 0 to q 20? c. At what rate is profit changing when q 20 recorders are produced? Is the profit increasing or decreasing at this level of production?

alones (alones):

Hrmm for each one?!? x.x ANywyas for A: \(\ P = qp - $70q = q ~(p - $70) = q ~(120 - q - $70) = 50q - q^2\) \[\ B.\] \(\ DP/dq = 50 - 2q \) \[\ C.\] \(\ dP/dq(20)\) Which would equal \(\ = $10\)